OpenStudy (anonymous):

Verify the idenity: 1-cos2x/1+sinx=sin x

OpenStudy (anonymous):

1 - cos2x = 2sin^2(x)

OpenStudy (anonymous):

is a rule. should help. lmk if i should solve

OpenStudy (anonymous):

you're making it more complicated :P

OpenStudy (dan815):

ya i think so lol

OpenStudy (anonymous):

2sinx = 1 + sinx can be obtained from the first line

OpenStudy (anonymous):

by dividing both sides by sinx

OpenStudy (dan815):

ah

OpenStudy (anonymous):

answer is sinx = 1/2 --> x = pi/6

OpenStudy (dan815):

hmm? its just a proof

OpenStudy (dumbcow):

\[2\sin^{2} x = \sin x + \sin^{2} x\] \[\sin^{2} x = \sin x\] \[\sin^{2} x - \sin x = 0\] \[\sin x (\sin x - 1) = 0\] x = 0, pi/2,pi

OpenStudy (dan815):

why do u solve for x

OpenStudy (dumbcow):

hmm its not an identity if its not true for all "x"

OpenStudy (dan815):

oh i see okay

OpenStudy (anonymous):

You need to prove how the left hand side equals to sin x when simplified.

OpenStudy (dan815):

its not possible in this case, thats what i thought

OpenStudy (dan815):

its asking u to verify if it equal

OpenStudy (dumbcow):

maybe its not written clearly...can you use the equation editor

OpenStudy (anonymous):

Verify each identity: \[1+\frac{ \cos ^{2}x }{ 1+\sin x }=\sin x\]

OpenStudy (anonymous):

1-

OpenStudy (dumbcow):

ok \[\cos^{2} x = 1-\sin^{2} x = (1+\sin x)(1-\sin x)\]

OpenStudy (anonymous):

\[1-\frac{(1+sinx)(1-sinx) }{1+\sin x }=sinx\] \[1- (1-sinx)=\sin x\]

OpenStudy (anonymous):

From here what happens to the negative infront of the sin x so that it can be equal to positive sin x

OpenStudy (dumbcow):

2 neg make a pos -(-sin x) = sin x